For the reaction NH4Cl(aq) NHj(g) + HCl(aq) ??? = 86.4 kJ and ? S°-79.1 J/K AG° would be negative at temperatures (above, below) Enter above or below in the first box and enter the temperature in the second box. Assume that AH° and ASo are constant. K. For the reaction Fe2O3(s) +3H2(g *2Fe(s) + 3i^0(g) ??? = 98.8 kJ and ? S°-143 J/K AG° for this reaction would be negative at temperatures (above, below) Enter above or below in the first box and enter the temperature in the second box. Assume that ?11° and ? K. are constant.
Answer #1

1) DG0 = DH0-TDS0

at equilibrium,DG = 0

0 = 86.4-(x*79.1*10^-3)

x = T = 1092.3 k

above 1092.3 k ,DG0 will be -ve.(spontaneous)


2)

DG0 = DH0-TDS0

at equilibrium,DG = 0

0 = 98.8-(x*143*10^-3)

x = T = 690.91 k

above 690.91 k , DG0 will be -ve.(spontaneous)